# DistPred: A Distribution-Free Probabilistic Inference Method for Regression and Forecasting

### v1

@@ Abstract
Traditional regression and prediction tasks often only provide deterministic point estimates. To estimate the uncertainty or distribution information of the response variable, methods such as Bayesian inference, model ensembling, or MC Dropout are typically used. These methods either assume that the posterior distribution of samples follows a Gaussian process or require thousands of forward passes for sample generation. We propose a novel approach called DistPred for regression and forecasting tasks, which overcomes the limitations of existing methods while remaining simple and powerful. Specifically, we transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. We have compared our method with several existing approaches on multiple datasets and achieved state-of-the-art performance. Additionally, our method significantly improves computational efficiency. For example, compared to state-of-the-art models, DistPred has a 90x faster inference speed. Experimental results can be reproduced through this Repository.
@@ 1 Introduction
In this paper, we consider the underlying distribution behind predicting the response variable because it reflects the confidence intervals of all levels. For example, based on this distribution, we can calculate confidence intervals at any level, coverage rate, and uncertainty quantification. Currently, predicting the distribution of the response variable poses a challenge because at a specific moment, the response variable can only take on a single deterministic value. This point can be viewed as a maximum likelihood sample from its underlying distribution, but it fails to reflect the overall state of the underlying distribution.
Currently, the primary method used to address distribution prediction and uncertainty quantification in regression and prediction tasks is frequentist sampling. These methods involve sampling numerous samples by perturbing the explanatory variable or model to approximate the underlying distribution of the response variable. For instance, Bayesian neural networks (BNNs) simulate this uncertainty by assuming that their parameters follow a Gaussian distribution, thereby capturing the model’s uncertainty given the data (Blundell et al., 2015). Similarly, ensemble-based methods have been proposed to combine multiple deep models with random outputs to capture prediction uncertainty. MC Dropout (Gal and Ghahramani, 2016) demonstrates that enabling dropout during each testing process yields results akin to model ensembling. Additionally, models based on GANs and diffusion are introduced for conditional density estimation and prediction uncertainty quantification. These models utilize noise during the generation or diffusion process to obtain different predicted values for estimating the uncertainty of the response variable.
The common characteristic of these methods mentioned above is the requirement of $K$ forward passes to sample $K$ representative samples. For example, Bayesian framework-based methods require $K$ learnable parameter samples to be inferred to obtain $K$ representative samples; ensemble methods require $K$ models to jointly infer; MC Dropout requires $K$ forward passes with random dropout activation; generative models require $K$ forward or diffusion processes. However, the excessive forward passes result in significant computational overhead and slow speed, a drawback that becomes increasingly apparent for AI applications with high real-time requirements.
Furthermore, several methods (Nix and Weigend, 1994; Salinas et al., 2020; He et al., 2020) transform distribution prediction and uncertainty quantification into predicting statistical variables such as mean and variance by assuming that the response variable follows a known continuous distribution. For instance, DeepAR (Salinas et al., 2020) assumes the response variable follows a Gaussian distribution, thereby directly utilizing GaussianNLLLoss (Nix and Weigend, 1994) to optimize its mean and variance. These methods only predict statistical variables, reducing the inference cost, but strong distribution assumptions often fail to capture the true data distribution, resulting in inferior performance.
Additionally, there are methods that train models using quantiles as loss functions to obtain specific quantile intervals (QIs) of the response variable. For example, to predict the 5th and 95th percentile confidence intervals, one can directly predict the corresponding quantiles. MRQNN predicts specified quantiles using quantile loss. Similarly, TFT enables the model to directly output three quantiles: 0.1, 0.5, and 0.9. These methods cannot provide distribution information about the response variable but only offer fixed QIs, thus lacking flexibility in their predictions.
@@
Figure 1: DistPred can provide $K$ predicted values $\hat{y}$ of the response variable $y$ given the predictor variable $x$ in a single forward process, denoted as $\mathbb{E}(\hat{Y}|x)$, where $\hat{Y}$ represents a maximum likelihood sample of $y$. Based on this sampling, the probability mass distribution (PMD) $P_{\theta}(\hat{y}|x)$, cumulative distribution function (PDF) $F_{\theta}(\hat{y}|x)$, and confidence curve $CC_{\theta}(\hat{y}|x)$ for the response variable $y$ can be computed, thereby yielding comprehensive statistical insights into $y$. For instance, this includes confidence intervals (CI) at any desired level, as well as p-values.
@@
To address this issue, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks. DistPred is a simple and powerful method that can estimate the distribution of the response variable in a single forward pass. Specifically, we contemplate employing all predictive quantiles to specify the potential cumulative density function (CDF) of the predictor variable, and we show that the full quantiles’ prediction can be translated into calculating the minimum expected score of the response variable and the predictive ensemble variables. Based on this, we transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance. Further, we show that DistPred can provide comprehensive statistical insights into the response variable, including confidence intervals at any desired level, p-values, and other statistical information, as shown in Figure 1. Experimental results show that DistPred outperforms existing methods in terms of both accuracy and computational efficiency. Specifically, DistPred has a 90x faster inference speed than state-of-the-art models.
@@ 2 Method
Assume that the dataset $D=\{x_{i},y_{i}\}_{i=1}^{N}$ consists of $N$ sample-label pairs. The subscript $i$ will be omitted if it does not cause ambiguity in the context. Our objective is to utilize a machine learning model $M$ with parameters $\theta$ to predict the underlying distribution $P(y)$ of the response variable $y$ from $D$, aiming to acquire comprehensive statistical insights such as obtaining confidence intervals (CI) and quantifying uncertainty at any desired level.
Direct prediction of distribution $P$ is not feasible, because:
@@
    Without distributional assumptions, we cannot give a valid representation of the PDF or CDF of the predicted distribution $P_{\theta}(\hat{y})$.
@@
    We can only obtain a singular deterministic value for the response variable $y$, lacking access to its distributional information to guide model learning.
Figure 2: Relationship between CDF and all predictive quantiles.
To address the aforementioned issues, we contemplate employing all predictive quantiles $\hat{q}_{1},\hat{q}_{2},\cdots$, at levels $\alpha_{1},\alpha_{2},\cdots$ to specify the potential CDF $F_{\theta}(\hat{y})$ of the predictor variable $\hat{y}$. This is because if we know the cumulative distribution function of a random variable, we can find any quantile by setting $F(y)=q$. Conversely, if we have a complete set of quantiles, we can approximate or reconstruct the cumulative distribution function of the random variable. As shwon in Figure 2, quantiles provide discrete ‘snapshots’ of the distribution, while the CDF is a continuous, smooth version of these snapshots, offering a complete description of the cumulative probability from the minimum to the maximum value.
Next, we will present that the full predictive quantiles outlined earlier serves as a proper approximation of the response variable’s CDF. And the full quantiles’ prediction can be translated into calculating the minimum expected score of the response variable and the predictive ensemble variables $\hat{Y}$, where $\hat{Y}=\{\hat{y}_{1},\cdots,\hat{y}_{K}\}$. Prior to delving into this analysis, we shall initially introduce the scoring rules utilized for assessing the propriety of predicted distributions.
2.1 Utilize proper scoring rules as the loss function
Scoring rules offer a concise measure for assessing probabilistic forecasts by assigning numerical scores according to the forecast distribution and predicted outcomes (Gneiting and Raftery, 2007; Jordan et al., 2017). Specifically, Let $\Omega$ denote the set of possible values of the quantity of interest, and let $\mathcal{P}$ denote a convex class of probability distributions on $\Omega$. A scoring rule is a function
@@ 2.1 Utilize proper scoring rules as the loss function
for all $P,Q\in\mathcal{P}$. A scoring rule is termed strictly proper when equality is achieved only when $P=Q$. Proper scoring rules (PSR) are essential for comparative evaluation, particularly in ranking forecasts. In practice, the forecaster with the lowest average score across multiple forecast cases typically demonstrates the best predictive performance. Proper scoring rules incentivize forecasters to accurately report their perception of the true distribution in this scenario. Therefore, PSR provide attractive loss and utility functions that can be tailored to a regression or forecast problem. To estimate $\theta$, we might measure the goodness-of-fit by the mean score
$S_{n}(\theta)=\frac{1}{N}\sum_{i=1}^{N}S(P_{\theta}(\hat{y}_{i}),y_{i}).$ (3)
Let $\theta^{*}$ denotes the true parameter value, then asymptotic arguments indicate that $\text{argmin}_{\theta}S_{n}(\theta)\rightarrow\theta^{*}$ as $n\rightarrow\infty$. This suggests a general approach to transform PSR into loss functions for training models, which implicitly minimizes the divergence between predictive and true distributions.
2.2 Property of full predictive quantiles
We contemplate probabilistic predictions pertaining to a continuous quantity, manifested as full predictive quantiles $\hat{q}_{1},\cdots,\hat{q}_{K}$. For $P\in\mathcal{P}$, let $q_{1},\cdots,q_{K}$ denote the true $P$-quantiles at levels $\alpha_{1},\cdots,\alpha_{K}\in(0,1)$. Then, the expected score $S(q_{1},\cdots,q_{K};P)$ can be defined as
Further, a scoring rule S is proper if $S(q_{1},\cdots,q_{K};P)\geq S(\hat{q}_{1},\cdots,\hat{q}_{K};P)$. Based on this definition, we assuming that $s_{k},i\in[1,\cdots K]$ is nondecreasing and $h$ is arbitrary, then the scoring rule
$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\sum_{k=1}^{K}\left(\alpha_{i}s_{k}(\hat{q}_{k})+(s_{k}(y)-s_{k}(\hat{q_{k}})\mathbbm{1}\{y\leq\hat{q}_{k}\})\right)$ (5)
is proper for predicting the quantiles at levels $\alpha_{1},\cdots,\alpha_{K}$ when $K\rightarrow\infty$ (Gneiting and Raftery, 2007; Schervish et al., 2012). $\mathbbm{1}\{y\leq\hat{q}_{k}\}$ denotes the indicator function which is one if $y\leq\hat{q}_{k}$ and zero otherwise.
Equation 5 shows that full predictive quantiles is proper. In essence, delineating a predictive CDF is equivalent to specifying all predictive quantiles. Consequently, we can formulate scoring rules for the predictive distribution based on the scoring rules for the quantiles. Specifically, let $S_{\alpha}$ denotes a proper scoring rule for the quantile at level $\alpha$, then the scoring rule
$S(F,y)=\int_{0}^{1}S_{\alpha}(F^{-1}(\alpha);y)\,\text{d}\alpha=\int_{-\infty}^{\infty}S(F(\hat{y}),\mathbbm{1}\{y\leq\hat{y}\})\,\text{d}\hat{y}$ (6)
is proper. Here, we can find that the right of Equation 6 corresponds to the CRPS in which $S$ is the quadratic or Brier score, which is defined as
$C(F,y)=\int_{-\infty}^{\infty}(F(\hat{y})-\mathbbm{1}\{y\leq\hat{y}\})^{2}\,\text{d}\hat{y}.$ (7)
If the first moment of $F$ is finite, the CRPS can be written as
@@ 2.2 Property of full predictive quantiles
where $\hat{Y}$ and $\hat{Y}^{\prime}$ denote independent predictive variables with distribution $F$.
Figure 3: The workflow of DistPred. An ensemble of predictive variables $\hat{Y}$ is inferred in a forward pass and PSR $S(\mathbb{E}(\hat{Y}|x),y)$ is utilized to train the learner end-to-end.
@@ 2.3 End-to-end ensemble inference
Figure 4: DistPred’s architecture.
Based on the analysis provided above, it is evident that predicting the full quantiles is equivalent to minimizing Equation 8 w.r.t $\mathbb{E}(\hat{Y}|y)$. Hence, as the workflow shown in Figure 3, we can develop a model $M$ with parameters $\theta$ that infers an ensemble of predictive variables $\hat{Y}$ in a forward pass and utilize Equation 8 to train it end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the empirical CDF $\hat{F}$ by the predictive ensemble variables
$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|-\frac{1}{2K^{2}}\sum_{k=1}^{K}\sum_{j=1}^{K}|\hat{y}_{k}-\hat{y}^{\prime}_{j}|.$ (9)
It’s worth noting that Equation 9 is a differentiable discrete form w.r.t $\hat{Y}$ and $\hat{Y^{\prime}}$ that strictly satisfies PSR. However, implementations of Equation 9 exhibit inefficiency due to their computational complexity of $O(K^{2})$. This can be enhanced by employing representations based on the generalized quantile function (Laio and Tamea, 2007) and the sorted predictive ensemble variables
$C(\hat{F},y)=\frac{1}{2K^{2}}\sum_{k=1}^{K}(\overrightarrow{y_{k}}-y)(k\mathbbm{1}\{y\leq\overrightarrow{y_{k}}\}-i+\frac{1}{2}).$ (10)
Since a sorting operation is involved, the computational complexity of Equation 10 is $O(KlogK)$. To conserve memory, we also suggest utilizing Equation 10 as the loss function, as predicting ensemble variables in long-term forecasting tasks may lead to out-of-memory issues on GPUs.

### v2

@@ Abstract
Traditional regression and prediction tasks often only provide deterministic point estimates. To estimate the distribution or uncertainty of the response variable, traditional methods either assume that the posterior distribution of samples follows a Gaussian process or require thousands of forward passes for sample generation. We propose a novel approach called DistPred for regression and forecasting tasks, which overcomes the limitations of existing methods while remaining simple and powerful. Specifically, we transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. We have compared our method with several existing approaches on multiple datasets and achieved state-of-the-art performance. Additionally, our method significantly improves computational efficiency. For example, compared to state-of-the-art models, DistPred has a 180x faster inference speed. Experimental results can be reproduced through this Repository.
@@ 1 Introduction
In this paper, we consider the underlying distribution behind predicting the response variable because it reflects the confidence intervals at all levels. For example, based on this distribution, we can calculate confidence intervals, coverage rate, and uncertainty quantification at any level, as shown in Fig. 1. Currently, predicting the distribution of the response variable poses a challenge because at a specific moment, the response variable can only take on a single deterministic value. This point can be viewed as a ‘representative’ sample from its underlying distribution, but it fails to represent the overall state of the underlying distribution.
@@
Figure 1: DistPred can provide $K$ predicted values $\hat{y}$ of the response variable $y$ given the predictor variable $x$ in a single forward process, denoted as $\mathbb{E}(\hat{Y}|x)$, where $\hat{Y}$ represents a maximum likelihood sample of $y$. Based on this sampling, the probability mass function (PMF) $P_{\theta}(\hat{y}|x)$, cumulative distribution function (CDF) $F_{\theta}(\hat{y}|x)$, and confidence curve $CC_{\theta}(\hat{y}|x)$ for the response variable $y$ can be computed, thereby yielding comprehensive statistical insights into $y$. For instance, this includes confidence intervals (CI) at any desired level, as well as p-values.
Currently, several methods are employed to predict the underlying distribution of the response variable in regression and forecasting tasks. The most straightforward approach is to assume that the response variable follows a prior distribution and represents a specific statistic from that distribution. Specifically, several methods (Bishop, 1994; Greene, 2003; Salinas et al., 2020; He et al., 2020) transform distribution prediction and uncertainty quantification into predicting statistical variables such as mean and variance by assuming that the response variable follows a known continuous distribution. For instance, mixture density networks (Bishop, 1994) superimpose a specific distribution, typically Gaussian, weighted by designated parameters to fit the prior distribution. Heteroscedasticity regression (Greene, 2003) quantifies uncertainty by modeling the variability of residuals as a function of independent variables. These methods only predict statistical variables, which reduces the inference cost, but strong assumptions often fail to capture the true data distribution, resulting in inferior performance.
Conformal prediction offers an alternative approach for distribution prediction (Vovk et al., 2017, 2018; Romano et al., 2019; Xu and Xie, 2021, 2023). The authors defined the random prediction system in (Vovk et al., 2017, 2018) and proposed a nonparametric prediction distribution method based on conformal assumptions. The authors integrated conformal prediction with quantile regression in (Romano et al., 2019; Xu and Xie, 2021, 2023) to construct prediction intervals for the response variable by training multiple bootstrap estimators. However, the application of conformal prediction to distribution prediction can be constrained by its reliance on the exchangeability assumption of residuals and the challenges in managing temporal dependencies (autocorrelation), potentially resulting in less reliable prediction intervals in non-i.i.d. data settings.
Uncertainty quantification is a method that indirectly reflects the potential distribution of the response variable, which can be classified into two primary categories: epistemic uncertainty and aleatoric uncertainty Der Kiureghian and Ditlevsen (2009). Epistemic uncertainty refers to the uncertainty within the model itself. In contrast, aleatoric uncertainty pertains to the inherent randomness in the observations. Quantitative analysis involves generating numerous samples (e.g., using MCMC) by perturbing the explanatory variables or models, thereby approximating the underlying distribution. For instance, bayesian neural networks (BNNs) (Blundell et al., 2015; Immer et al., 2021; Daxberger et al., 2021) simulate this uncertainty by assuming that their parameters follow a predefined distribution, thereby capturing the model’s uncertainty given the data. Similarly, ensemble-based methods have been proposed to combine multiple deep models with random outputs to capture prediction uncertainty. MC Dropout (Gal and Ghahramani, 2016) shows that enabling dropout during each testing process yields results akin to model ensembling. Additionally, models based on GANs and diffusion have been introduced for conditional density estimation and prediction uncertainty quantification (Zhou et al., 2021b; Liu et al., 2021; Han et al., 2022). These models utilize noise during the generation or diffusion process to obtain different predicted values for estimating the uncertainty of the response variable.
The common characteristic of these methods mentioned above is the requirement of $K$ forward passes to sample $K$ representative samples. For example, Bayesian framework-based methods require $K$ learnable parameter samples to be inferred in order to obtain $K$ representative samples; ensemble methods require $K$ models to jointly infer; MC Dropout requires $K$ forward passes with random dropout activations; generative models require $K$ forward or diffusion processes. However, the excessive forward passes result in significant computational overhead and slow speed, a drawback that becomes increasingly apparent for AI applications with high real-time requirements.
@@
To address this issue, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks. DistPred is a simple and powerful method that can estimate the distribution of the response variable in a single forward pass. Specifically, we contemplate employing all predictive quantiles to specify the potential cumulative density function (CDF) of the predictor variable, and we show that the full quantiles’ prediction can be translated into calculating the minimum expected score of the response variable and the predictive ensemble variables. Based on this, we transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance. Further, we show that DistPred can provide comprehensive statistical insights into the response variable, including confidence intervals at any desired level, p-values, and other statistical information, as shown in Fig. 1. Experimental results show that DistPred outperforms existing methods in terms of both accuracy and computational efficiency. Specifically, DistPred has a 180x faster inference speed than state-of-the-art models.
Figure 2: The workflow of DistPred. An ensemble of predictive variables $\hat{Y}$ is inferred in a forward pass and $S(\mathbb{E}(\hat{Y}|x),y)$ is utilized to train the learner end-to-end.
@@ 2 Method
Assume that the dataset $D=\{x_{i},y_{i}\}_{i=1}^{N}$ consists of $N$ sample-label pairs. The subscript $i$ will be omitted if it does not cause ambiguity in the context. Our objective is to utilize a machine learning model $M$ with parameters $\theta$ to predict the underlying distribution $P(y|x)$ from $D$, aiming to acquire comprehensive statistical insights such as obtaining confidence intervals (CI) and quantifying uncertainty at any desired level.
Direct prediction of distribution $P_{\theta}(\hat{y}|x)$ is not feasible because:
@@
    Without distributional assumptions, we cannot give a valid representation of the PDF or CDF for the response variable.
@@
    We have only a deterministic target point, without access to its distribution information, which limits our ability to guide the model’s learning process.
To address the aforementioned issues, we contemplate employing full predictive quantiles $\hat{q}_{1},\hat{q}_{2},\cdots,\hat{q}_{K}$ at levels $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$, $(K\rightarrow\infty)$, to specify the potential CDF $F_{\theta}(\hat{y})$ of the response variable $\hat{y}$. This is because if we know the cumulative distribution function of a random variable, we can find any quantile by
$q_{\alpha}=\text{inf}\{y\in\mathbb{R}:F(y)\geq\alpha\}.$ (1)
Conversely, if we have a complete set of quantiles, we can approximate or reconstruct the cumulative distribution function of the random variable
$F(y)=\text{sup}\{\alpha\in[0,1]:q_{\alpha}\leq y\}.$ (2)
Full quantiles provide discrete ‘snapshots’ of the distribution, while the CDF is a continuous, smooth version of these snapshots, offering a complete description of the cumulative probability from the minimum to the maximum value.
2.1 Property of Full Quantiles
We contemplate probabilistic predictions pertaining to a continuous quantity, manifested as full predictive quantiles $\hat{q}_{1},\cdots,\hat{q}_{K},(K\rightarrow\infty)$. For $P\in\mathcal{P}$, let $\hat{q}_{1},\cdots,\hat{q}_{K}$ denote the true $P$-quantiles at levels $\alpha_{1},\cdots,\alpha_{K}\in(0,1)$. Then, the expected score $S(q_{1},\cdots,q_{K};P)$ can be defined as
The function $S$ in Eq. 3 satisfies the scoring rule, which offers a concise measure for assessing probabilistic forecasts by assigning numerical scores according to the forecast distribution and predicted outcomes (Gneiting and Raftery, 2007; Jordan et al., 2017). Specifically, let $\Omega$ denote the set of possible values of the quantity of interest, and let $\mathcal{P}$ denote a convex class of probability distributions on $\Omega$, the scoring rule is a function
@@ 2.1 Property of Full Quantiles
for all $P,Q\in\mathcal{P}$. A scoring rule is termed strictly proper when equality is achieved only when $P=Q$. Proper scoring rules are essential for comparative evaluation, especially in ranking forecasts.
Based on this definition, we can get the scoring rule $S$ is proper when $S(q_{1},\cdots,q_{K};P)\geq S(\hat{q}_{1},\cdots,\hat{q}_{K};P)$. Further, assuming that $s_{k}$ is a non-decreasing probability measure of $\mathcal{P}$, then the scoring rule
$S(\hat{q}_{1},\cdots,\hat{q}_{K};P)=\sum_{k=1}^{K}\left(\alpha_{k}s_{k}(\hat{q}_{k})+(s_{k}(y)-s_{k}(\hat{q_{k}})\mathbb{I}\{y\leq\hat{q}_{k}\})\right)$ (6)
is proper for predicting the quantiles at levels $\alpha_{1},\cdots,\alpha_{K}$ when $K\rightarrow\infty$ (Gneiting and Raftery, 2007; Schervish et al., 2012). $\mathbb{I}\{y\leq\hat{q}_{k}\}$ denotes the indicator function which is $1$ if $y\leq\hat{q}_{k}$ and 00 otherwise.
2.2 Predicting CDF by Full Quantiles
Eq. 6 shows that full predictive quantiles are proper. Consequently, we can formulate scoring rules for the predictive distribution based on the scoring rules for the quantiles. Specifically, let $S_{\alpha}$ denote a proper scoring rule for the quantile at level $\alpha$, then the scoring rule
$S(F,y)=\int_{0}^{1}S_{\alpha}(F^{-1}(\alpha);y)\,\text{d}\alpha=\int_{-\infty}^{\infty}S(F(\hat{y}),\mathbb{I}\{y\leq\hat{y}\})\,\text{d}\hat{y}$ (7)
is proper. Here, we establish a relationship between the full quantiles of the response variable and its CDF. However, the CDF here is in continuous form and cannot be directly observed. Therefore, we need to convert it into a discrete form, as the number of quantiles $K$ is always finite in practice. It is worth noting that Eq. 7 corresponds to the continuous ranked probability score (CRPS) in which $S$ is the quadratic or Brier score, defined as
$C(F,y)=\int_{-\infty}^{\infty}(F(\hat{y})-\mathbb{I}\{y\leq\hat{y}\})^{2}\,\text{d}\hat{y}.$ (8)
When the first moment of $F$ is finite, the CRPS can be written as
@@ 2.2 Predicting CDF by Full Quantiles
where $\hat{Y}$ and $\hat{Y}^{\prime}$ denote independent predictive ensemble variables with distribution $F$.
Eq. 9 incentivizes forecasters to accurately report their perception of the true distribution in this scenario. Therefore, it provides attractive measures and utility functions that can be tailored to a regression or forecast problem. To estimate $\theta$, we need to measure the goodness-of-fit by the mean score
$C_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}C(F_{\theta}(\hat{y}_{i}),y_{i}).$ (10)
Let $\theta^{*}$ denote the true parameter value, then asymptotic arguments indicate that $\text{argmin}_{\theta}C_{N}(\theta)\rightarrow\theta^{*}$ as $N\rightarrow\infty$. As shown in Fig. 2, the workflow suggests a general approach to transforming proper scoring rules into loss functions for training models, which implicitly minimizes the discrepancy between predictive and true distributions.
@@ 2.3 End-to-End Ensemble Inference
Figure 3: DistPred’s architecture. A model $M$ with parameters $\theta$ takes one input variable $x$ and outputs an ensemble of $K$ response variable $\hat{y}_{1},\cdots,\hat{y}_{K}$.
Based on the analysis provided above, it is evident that predicting the full quantiles is equivalent to minimizing Eq. 9 w.r.t $\mathbb{E}(\hat{Y}|y)$. Hence, as illustrated in the architecture depicted in Fig. 3, we can develop a model $M$ with parameters $\theta$ that infers an ensemble of predictive variables $\hat{Y}=\{\hat{y}_{1},\cdots,\hat{y}_{K}\}$ in a forward pass and utilize Eq. 10 to train it end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the empirical CDF $\hat{F}$ by the predictive ensemble variables
$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|-\frac{1}{2K^{2}}\sum_{k=1}^{K}\sum_{j=1}^{K}|\hat{y}_{k}-\hat{y}_{j}|.$ (11)
It’s worth noting that Eq. 11 is a differentiable discrete form w.r.t $\hat{Y}$ and $\hat{Y^{\prime}}$ that strictly satisfies proper scoring rules. However, implementations of Eq. 11 exhibit inefficiency due to their storage complexity of $\mathcal{O}(K^{2})$. This can be enhanced by using algebraically equivalent representations based on the generalized quantile function (Laio and Tamea, 2007) and the sorted predictive ensemble variables $\overrightarrow{y_{k}}$,
$C(\hat{F},y)=\frac{1}{2K^{2}}\sum_{k=1}^{K}(\overrightarrow{y_{k}}-y)(k\mathbb{I}\{y\leq\overrightarrow{y_{k}}\}-k+\frac{1}{2}).$ (12)
Since a sorting operation is involved, the storage complexity of Eq. 12 is $\mathcal{O}(KlogK)$. However, since Eq. 12 renders the objective $C(\hat{F},y)$ non-differentiable, we employ probability-weighted moment estimation form to approximate it, as
$C(\hat{F},y)=\frac{1}{K}\sum_{k=1}^{K}|\hat{y}_{k}-y|+\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}-\frac{2}{K(K-1)}\sum_{k=1}^{K}\hat{y}_{k}(k-1).$ (13)
Actually, according to the definition of L-Moments, we can prove that the last term of Eq. 11 can be decomposed into the 1-th and 0-st order L-moments of Eq. 13, i.e., Eq. 13 is an unbiased estimate of Eq. 11 when it has a finite first L-moment. To conserve memory, we suggest utilizing Eq. 13, which has a storage complexity of $\mathcal{O}(K)$, as the loss function, since predicting ensemble variables in long-term forecasting tasks may lead to out-of-memory issues on GPUs.